A sample of 100 one-dollar bills from the Subway cash register revealed that 16 had something written on them besides the normal printing (e.g., “Bob タ Mary”). (a) At α = .05, is this sample evidence consistent with the hypothesis that 10 percent or fewer of all dollar bills have anything written on them besides the normal printing? (a) At α = .05, is this sample evidence consistent with the hypothesis that 10 percent or fewer of all dollar bills have anything written on them besides the normal printing?
Answer by stanbon(75887)

A sample of 100 one-dollar bills from the Subway cash register revealed that 16 had something written on them besides the normal printing (e.g., “Bob タ Mary”). (a) At α = .05, is this sample evidence consistent with the hypothesis that 10 percent or fewer of all dollar bills have anything written on them besides the normal printing?
-----------------
Ho: p = 0.10
Ha: p is not 0.10
---------------------
n = 100; p-hat = 16/100 = 0.16; alpha = 5%;
Critical values for 2-tail z-test: +/-1.96
-----------------
Test Statistic:
z(0.16) = (0.16-0.10)/sqrt[0.1*0.9/100] = 0.16/[0.0018] = 5.333..
-------------
Conclusion: Since the test statistic is in the reject interval,
reject Ho.
Conclude that p not not = 0.10
=================================
Cheers,
Stan H.